OpenStudy (anonymous):

Complex numbers~ How would simplify 4i^5

OpenStudy (ivettef365):

Do you know the rules of i^2, i^3, i^4

OpenStudy (anonymous):

Not Quite well

OpenStudy (ivettef365):

well i=i , i^2 = -1 , i^3 = -i and i^4=1

OpenStudy (ivettef365):

Knowing that you have 4i^5 then that is the same as (4)(i^4)(i)

OpenStudy (ivettef365):

now just plug in (4)(1)(i) that is equal to 4i

OpenStudy (anonymous):

(4)(1^4)(4i)

OpenStudy (ivettef365):

Is that another problem?

OpenStudy (anonymous):

lol no, im confused now was 4i the asnwer ?

OpenStudy (ivettef365):

Correct

OpenStudy (anonymous):

ok

OpenStudy (ivettef365):

4 x 1 = 4 and 4 x i = 4i

OpenStudy (anonymous):

\[4\iota ^{5}=4\left( \iota ^{2} \right)^{2}\iota=4\left( -1 \right)^{2}\iota=4\iota \]